# SLiM: One-shot Quantization and Sparsity with Low-rank Approximation for LLM Weight Compression

### v2

@@ Abstract
Conventional model compression techniques for LLMs address high memory consumption and slow inference challenges but typically require computationally expensive retraining to preserve accuracy. In contrast, one-shot compression methods eliminate retraining cost, but struggle to achieve accuracy comparable to dense models. This paper presents SLiM, a new one-shot compression framework that holistically integrates hardware-friendly quantization, sparsity, and low-rank approximation into a unified process. First, we formulate the quantization process using a probabilistic approach (SLiM-Quant) that enables us to apply uniform quantization. Then, we use an existing one-shot pruning method to apply semi-structured sparsity on top of the quantized weights. Finally, to compensate for the introduced aggregated quantization and sparsity error, we use a novel saliency function with unique invertible and additive features that enables us to mathematically compute the value of low-rank adapters. SLiM improves model accuracy by up to 5.66%  (LLaMA-2-7B) for 2:4 sparsity with 4-bit weight quantization, outperforming prior methods. Models compressed with SLiM achieve up to 3.78$\times$  and 3.75$\times$  layer-wise speedup on Nvidia RTX3060 and A100 GPUs, respectively. We also propose an optional PEFT recipe that further improves accuracy by up to 1.66%  (LLaMA-2-13B) compared to SLiM without fine-tuning. [^1] Code and data for SLiM is available at: https://github.com/Mohammad-Mozaffari/slim
@@ 1 Introduction
To eliminate the need for retraining, one-shot compression methods have gained traction, achieving high accuracy using only a relatively small set of calibration data. State-of-the-art methods such as SparseGPT (Frantar & Alistarh, 2023) and Wanda (Sun et al., 2023) have demonstrated strong one-shot pruning performance. However, these methods perform well with unstructured sparsity but struggle with semi-structured patterns, such as NVIDIA’s 2:4 sparsity pattern (Mishra et al., 2021), which is crucial for efficient hardware-accelerated inference. One-shot quantization methods use various methods including row/column reordering (Frantar & Alistarh, 2023), scaling (Xiao et al., 2023; Lin et al., 2024), and other transformations (Shao et al., 2023; Ma et al., 2024) to mitigate model accuracy loss. However, their dependence on complex GPU kernel implementations can limit the practical benefits of quantization and add challenges to their development.
@@
Low-rank adapters have emerged as a promising approach for mitigating the accuracy loss introduced by model compression techniques. Recent studies (Guo et al., 2023; Li et al., 2023) have explored learnable low-rank adapters to reduce weight reconstruction errors caused by pruning or quantization. However, these methods typically require an expensive retraining process on hundreds of millions of tokens to recover performance (Dettmers et al., 2023; Nikdan et al., 2024). This retraining is necessary because low-rank adapters are often initialized based on weight norms rather than their direct impact on model outputs, resulting in suboptimal starting points (Guo et al., 2023). To address this limitation, L2QER (Zhang et al., 2024) introduces a one-shot low-rank adapter approach that compensates for quantization loss by directly minimizing its impact on model outputs. While effective for quantization, L2QER struggles to maintain accuracy when combined with sparsity, underscoring the need for compression techniques that seamlessly integrate low-rank approximations with both sparsity and quantization. [^2] For a more detailed discussion of the related work, see Appendix Q.
@@
To address these limitations, we propose SLiM, a one-shot compression framework that seamlessly integrates hardware-friendly sparsity, quantization, and low-rank approximation to minimize accuracy degradation while maintaining computational efficiency. We decompose the primary objective of SLiMone-shot hardware-friendly sparsity and quantization with minimal accuracy loss–into three key sub-tasks. For quantization, we prioritize uniform quantization (e.g., one scale per tensor) due to its computational efficiency on commodity hardware and its simplified encoding/decoding process. However, standard uniform quantization often introduces significant quantization errors, particularly in tensors with a wide dynamic range or outliers (Gholami et al., 2022). These errors degrade model accuracy, making uniform quantization less attractive than more complex per-group quantization methods. Optimizing uniform quantization is inherently a non-convex problem, and recent approaches (Nagel et al., 2021) rely on grid search to find an optimal scaling factor. However, grid search can be both suboptimal and computationally expensive, often leading to degraded model quality. To address this, we introduce SLiM-Quant, a probabilistic formulation of the quantization process. Our approach reformulates the inherently non-convex quantization problem into a convex optimization problem, allowing for an efficient and tractable solution. This transformation significantly reduces uniform quantization error, achieving accuracy levels comparable to more complex group quantization methods while retaining the computational efficiency of uniform quantization (up to 6% speedup). After quantization, we apply Wanda (Sun et al., 2023), a state-of-the-art pruning method, to introduce different forms of unstructured and structured sparsity on the quantized weights.
@@
Compared to state-of-the-art methods, SLiM achieves an average accuracy improvement of 5.66% on LLaMA-2-7B, 3.89% on LLaMA-2-13B, and 0.60% on OPT-13B under 2:4 sparsity and 4-bit weight quantization. More importantly, SLiM shifts the Pareto frontier, delivering higher model accuracy at the same total bit budget compared to existing compression techniques (up to 0.5%) and even outperforming dense models (up to 0.6%). These results highlight the effectiveness of SLiM in maximizing model quality under stringent resource constraints, making it an appealing solution for efficient large-scale deployment. Beyond accuracy improvements, SLiM achieves up to 3.78$\times$  and 3.75$\times$  layer-wise speedup on NVIDIA RTX3060 and A100 GPUs, respectively, demonstrating its efficiency on modern hardware. Finally, to further narrow the gap between compressed and dense models, we use an optional lightweight PEFT method, which provides up to an 1.66%  additional accuracy improvement for LLaMA-2-13B.
@@ 2 Preliminaries
Group quantization (Alistarh et al., 2017; Gunho et al., 2022) tackles AbsMax’s outlier sensitivity by assigning separate scaling factors to subgroups of the weight matrix. This approach captures local variations in weight magnitudes, reducing quantization error for non-uniform distributions. However, storing multiple scaling factors increases memory usage, and subgroup-specific dequantization increase computational complexity, potentially slowing inference. The challenges of using group quantization are discussed in Appendix S.
@@ 3.2 SLiM-LoRA low-rank adapters
A straightforward approach minimizes the total error norm between $\mathcal{W}$ and $\mathcal{W}^{C}$, focusing solely on reducing the error magnitude while ignoring the saliency of individual elements in the weight matrix. We call this method Naive-LoRA as it overlooks the significance of individual elements in the weight matrix. However, this method is suboptimal and can be substantially improved.
@@
Assuming that there exists an additive invertible saliency function $F:{\mathbb{R}}^{d_{in}\times d_{out}}\rightarrow{\mathbb{R}}^{d_{in}\times d_{out}}$, we need to solve Equation 8 to find the optimal adapters. By using the additive property of the saliency function $F(.)$, we can simplify Equation 8 to Equation 9.
@@ 4 Experimental results
Models, Datasets, and Evaluation. [^3] All the experiments in the paper were run at the University of Toronto We evaluate SLiM on the OPT (Zhang et al., 2022) and LLaMA-2 (Touvron et al., 2023) model families, both of which serve as standard baselines in model compression studies (Ma et al., 2024; Frantar & Alistarh, 2023; Sun et al., 2023). Model accuracy is assessed on a range of zero-shot downstream tasks, including MMLU (Hendrycks et al., 2020), Piqa (Bisk et al., 2020), Arc-Easy, Arc-Challenge (Clark et al., 2018), WinoGrande (Sakaguchi et al., 2021), and OpenBookQA (Mihaylov et al., 2018). For zero-shot evaluations, we utilize the Language Model Evaluation Harness (Gao et al., 2024) framework. In line with prior work (Sun et al., 2023; Frantar & Alistarh, 2023; Ma et al., 2024), we also report the perplexity of the models on a language modeling task on the WikiText2 (Merity et al., 2016) dataset, provided in Appendix E.
@@
Baselines. We compare SLiM against state-of-the-art one-shot pruning methods, including Wanda (Sun et al., 2023), SparseGPT (Frantar & Alistarh, 2023), and Magnitude Pruning (Han et al., 2015), as well as one-shot quantization techniques like OPTQ (Frantar et al., 2022), OmniQuant (Shao et al., 2023), AffineQuant (Ma et al., 2024), L2QER (Zhang et al., 2024), and AbsMax. Additionally, we extend Joint Sparsification and Quantization (JSQ) (Guo et al., 2024) to support 4-bit weight quantization and include it in our experiments. To ensure fairness, we use the optimal hyperparameters reported for each method, or the default hyperparameters if not explicitly reported. For a thorough description of the notations used to show the different variants of SLiM, please see Table 4 in Appendix A. For more details about the hyperparameters used in different experiments, please see Appendix R.
@@
To the best of our knowledge, L2QER is the only compression method utilizing zero-shot low-rank adapters to enhance model accuracy. Our approach, SLiM, significantly diverges from L2QER in several key aspects. First, we employ saliency-based low-rank adapters to mitigate compression loss in quantized and sparse models, whereas L2QER is tailored exclusively for quantization, resulting in reduced accuracy when combined with sparsity, as demonstrated in the subsequent subsections. Second, we introduce SLiM-Quant , which lowers the overhead and complexity of group quantization compared to methods like L2QER. Finally, SLiM compresses and fine-tunes low-rank adapters efficiently to minimize overhead. In contrast, L2QER relies on full-precision low-rank adapters, which incur additional overhead and do not benefit from the parameter-efficient fine-tuning proposed in our work.
Accuracy results. We evaluate the accuracy of SLiMand other state-of-the-art pruning and quantization methods across 2:4 and unstructured sparsity benchmarks, highlighting SLiM’s superiority in Table 1. SparseGPT and Group OPTQ, designed to work together, achieve competitive performance. For other advanced quantization methods, we pruned models using Wanda and quantized the sparse checkpoints with Group AbsMax, AWQ, OmniQuant, and AffineQuant, reporting the best results (detailed in Appendix F). Notably, methods like OmniQuant and AffineQuant struggle to quantize OPT-350M, often resulting in NaN values. Moreover, AWQ, OmniQuant, AffineQuant, and L2QER encounter out-of-memory (OOM) errors when compressing models on a single A100-40GB GPU. While JSQ performs well for the LLaMA-2 family, its difficulty compressing the OPT family limits its broader applicability.
@@
Naive-LoRA SLiM-Quant 34.28 33.38 38.36 41.21 44.91 45.25 48.45 51.94
SLiM-LoRA SLiM-Quant 34.62 34.36 40.61 42.73 45.99 46.09 51.15 54.94
SLiM-LoRAQ SLiM-Quant 34.43 34.30 40.11 42.37 46.33 46.24 51.02 53.55
@@
Naive-LoRA SLiM-Quant 34.77 34.23 40.40 43.37 46.64 47.30 51.52 55.33
SLiM-LoRA SLiM-Quant 35.20 35.32 41.85 43.48 47.08 47.96 54.26 57.85
SLiM-LoRAQ SLiM-Quant 35.35 35.13 41.74 43.63 47.16 47.86 54.18 57.33
Table 2 demonstrates how lightweight fine-tuning (FT) improves the accuracy of both SLiM-LoRA and Naive-LoRA, with SLiM-LoRA exhibiting greater gains due to its saliency-aware design. Further details on the fine-tuning process and its overhead are provided in Appendix I, illustrating its practicality for enhancing compressed model performance.
@@
Naive-LoRA + FT SLiM-Quant 50.89 55.70
SLiM-LoRA + FT SLiM-Quant 52.12 56.60
SLiM-LoRAQ + FT SLiM-Quant 48.31 56.50
@@
Naive-LoRA + FT SLiM-Quant 52.90 57.08
SLiM-LoRA + FT SLiM-Quant 54.69 57.96
SLiM-LoRAQ + FT SLiM-Quant 53.57 57.78
Comparison of large compressed and small dense models This section compares large compressed models with dense models of equivalent parameter size, offering guidelines for configuration selection under hardware constraints. We focus on 2:4 sparsity due to its hardware acceleration support and evaluate the OPT model family, which spans a wide range of sizes for comprehensive analysis.
We analyze model performance by plotting average accuracy against parameter size, calculated as detailed in Appendix J. This visualization enables a direct performance comparison between models with an equal number of bits.
@@
Speedup. Leveraging sparsity and quantization enhances GPU resource utilization, enabling faster model inference. Following Wanda’s experimental setup, we evaluate the speedup achieved across different model layers and sizes. Similar to Wanda, AWQ, and QuaRot (Ashkboos et al., 2024), we focus on consumer-grade GPUs and conduct our experiments on NVIDIA RTX 3060 GPUs. Speedup results for NVIDIA A100 GPUs are provided in Appendix H.
SLiM achieves notable speedups through optimized sparse and quantized matrix multiplication, utilizing Sparse Marlin (Frantar et al., 2024) integrated with vLLM (Kwon et al., 2023). For inference, we adopt small batch sizes during decoding, as recommended by prior works (Xia et al., 2023; Zheng et al., 2022). Dense Quantized Marlin or PyTorch kernels handle the low-rank adapters based on their quantization status. Table 3 highlights the speedup achieved across different LLaMA-2 layers compared to dense, unquantized models. Larger matrices, such as those in self-attention and feed-forward modules, consistently yield greater speedups, aligning with trends detailed in Appendix H.
@@
We evaluate SLiM on sparse-only and quantized-only models to isolate their effectiveness. Results in Additional Sparse-only Results (Appendix B) and Additional Quantization-only Results (Appendix C) demonstrate that SLiM and SLiM-Quant consistently outperform state-of-the-art compression methods.
The Language Modeling Experiments (Appendix E) evaluate SLiM across sparse and quantized, sparse-only, and quantized-only models on WikiText-2. The results align with the accuracy trends reported in the main paper, further validating the effectiveness of SLiM .
The Fine-tuning Costs (Appendix I) show that SLiM reduces fine-tuning overhead from over 36 days for 13B parameter models to just 14 hours on a single GPU, demonstrating its practicality and efficiency.
Table 3: LLaMA-2 family of models speedup ($\times$) using SLiM compared to original dense unquantized model on NVIDIA RTX-3060. $\uparrow$ shows higher speedup.
Model Batch LoRA Self- Up- Down
Size Size Type Attention Projection Projection
7B 16 FP16 1.99 2.58 3.23
INT4 1.00 2.29 2.58
32 FP16 2.06 2.53 3.40
INT4 1.13 2.27 2.97
64 FP16 1.54 1.70 1.85
INT4 0.96 1.62 1.76
13B 16 FP16 2.18 2.53 2.60
INT4 1.28 3.24 3.17
32 FP16 2.23 2.68 2.91
INT4 1.43 2.96 3.20
64 FP16 1.38 1.78 1.67
INT4 1.21 1.69 1.65
70B 16 FP16 2.18 2.86 2.75
INT4 3.11 3.99 3.79
32 FP16 2.00 2.63 2.67
INT4 2.75 3.19 3.39
64 FP16 1.38 1.70 1.86
INT4 1.51 1.77 1.94
We provide a comparison between Sparsity vs. Quantization (Appendix G) to show that combining 50% sparsity and 4-bit quantization helps achieve better compression results in comparison to solely using 2-bit quantization, while maintaining a similar compression ratio ($\sim$8×8$\times$8 ×).
Additional speedup results for SLiM on NVIDIA A100-40GB GPUs are provided in the Additional Speedup Results (Appendix H). A theoretical analysis of computation and memory reductions can be found in the Computation Reduction Analysis (Appendix K) and Memory Reduction Analysis (Appendix J), highlighting the efficiency of SLiM .
Compression Costs (Appendix L) details the time required to compress models of various sizes across different methods. Rank Analysis (Appendix M) explores how rank choices in low-rank adapters impact computational and memory costs, as well as model accuracy. Sparsity Analysis (Appendix B) analyzes the effects of different sparsity ratios on model compression. Lastly, Effects of Calibration Sample Count (Appendix N) evaluates the influence of calibration sample counts on the accuracy of calibration-based methods.

### v3

@@ Abstract
Conventional model compression techniques for LLMs address high memory consumption and slow inference challenges but typically require computationally expensive retraining to preserve accuracy. In contrast, one-shot compression methods eliminate retraining costs, but struggle to achieve accuracy comparable to dense models. This paper presents SLiM, a new one-shot compression framework that holistically integrates hardware-friendly quantization, sparsity, and low-rank approximation into a unified process. First, we formulate the quantization process using a probabilistic approach (SLiM-Quant) that enables us to apply uniform quantization. Then, we use an existing one-shot pruning method to apply semi-structured sparsity on top of the quantized weights. Finally, to compensate for the introduced aggregated quantization and sparsity error, we use a novel saliency function with unique invertible and additive features that enables us to mathematically compute the value of low-rank adapters. SLiM improves model accuracy by up to 5.66%  (LLaMA-2-7B) for 2:4 sparsity with 4-bit weight quantization, outperforming prior methods. Models compressed with SLiM achieve up to 4.3$\times$  and 3.8$\times$  layer-wise speedup on Nvidia RTX3060 and A100 GPUs, respectively. Additionally, they achieve up to 0.23×0.23$\times$0.23 × end-to-end memory reduction in comparison to their dense counterparts. We also propose an optional PEFT recipe that further improves accuracy by up to 1.66%  (LLaMA-2-13B) compared to SLiM without fine-tuning. [^1] Code and data for SLiM is available at: https://github.com/Mohammad-Mozaffari/slim
@@ 1 Introduction
To eliminate the need for retraining, one-shot compression methods have gained traction, achieving high accuracy using only a relatively small set of calibration data. State-of-the-art methods such as SparseGPT (Frantar & Alistarh, 2023) and Wanda (Sun et al., 2023) have demonstrated strong one-shot pruning performance. However, these methods perform well with unstructured sparsity but struggle with semi-structured patterns, such as NVIDIA’s 2:4 sparsity pattern (Mishra et al., 2021), which is crucial for efficient hardware-accelerated inference. To solve this challenge, unlike SparseGPT and Wanda that focus on layer-wise error minimization, MaskLLM (Fang et al., 2024) uses a learnable mask that minimizes the end-to-end model error, resulting in a significant boost to the accuracy of the model at the cost of an expensive mask training phase. One-shot quantization methods use various methods including row/column reordering (Frantar & Alistarh, 2023), scaling (Xiao et al., 2023; Lin et al., 2024), and other transformations (Shao et al., 2023; Ma et al., 2024) to mitigate model accuracy loss. However, their dependence on complex GPU kernel implementations can limit the practical benefits of quantization and add challenges to their development.
@@
Low-rank adapters have emerged as a promising approach to mitigate the accuracy loss introduced by model compression techniques. Recent studies (Guo et al., 2023; Li et al., 2023) have explored learnable low-rank adapters to reduce weight reconstruction errors caused by pruning or quantization. However, these methods typically require an expensive retraining process on hundreds of millions of tokens to recover performance (Dettmers et al., 2023; Nikdan et al., 2024). This retraining is necessary because low-rank adapters are often initialized based on weight norms rather than their direct impact on model outputs, resulting in suboptimal starting points (Guo et al., 2023). To address this limitation, L2QER (Zhang et al., 2024a) introduces a one-shot low-rank adapter approach that compensates for quantization loss by directly minimizing its impact on model outputs. QERA (Zhang et al., 2024b) and CALDERA (Saha et al., 2024) find closed form solutions to the problem that L2QER tries to solve. Although effective for quantization, methods such as L2QER struggle to maintain accuracy when combined with sparsity, underscoring the need for compression techniques that seamlessly integrate low-rank approximations with both sparsity and quantization [^2] For a more detailed discussion of the related work, see Appendix S.
@@
To address these limitations, we propose SLiM, a one-shot compression framework that seamlessly integrates hardware-friendly sparsity, quantization, and low-rank approximation to minimize accuracy degradation while maintaining computational efficiency. We decompose the primary objective of SLiMone-shot hardware-friendly sparsity and quantization with minimal accuracy loss–into three key sub-tasks. For quantization, we prioritize uniform quantization (e.g., one scale per tensor) because of its computational efficiency on commodity hardware and its simplified encoding/decoding process. However, standard uniform quantization often introduces significant quantization errors, particularly in tensors with a wide dynamic range or outliers (Gholami et al., 2022). These errors degrade model accuracy, making uniform quantization less attractive than more complex per-group quantization methods. Optimizing uniform quantization is inherently a non-convex problem, and recent approaches (Nagel et al., 2021) rely on grid search to find an optimal scaling factor. However, grid search can be both suboptimal and computationally expensive, often leading to poor model quality. To address this, we introduce SLiM-Quant, a probabilistic formulation of the quantization process. Our approach reformulates the inherently non-convex quantization problem into a convex optimization problem, allowing for an efficient and tractable solution. This transformation significantly reduces uniform quantization error, achieving accuracy levels comparable to more complex group quantization methods while retaining the computational efficiency of uniform quantization (up to 6% speedup). After quantization, we apply Wanda (Sun et al., 2023), a state-of-the-art pruning method, to introduce different forms of unstructured and structured sparsity on the quantized weights.
@@
Compared to state-of-the-art methods, SLiM achieves an average accuracy improvement of 5.66% on LLaMA-2-7B, 3.89% on LLaMA-2-13B, and 0.60% on OPT-13B under 2:4 sparsity and 4-bit weight quantization. More importantly, SLiM shifts the Pareto frontier, delivering higher model accuracy at the same total bit budget compared to existing compression techniques (up to 0.5%) and even outperforming dense models (up to 0.6%). These results highlight the effectiveness of SLiM in maximizing model quality under stringent resource constraints, making it an appealing solution for efficient large-scale deployment. Beyond accuracy improvements, SLiM achieves up to 4.3$\times$  and 3.8$\times$  layer-wise speedup on NVIDIA RTX3060 and A100 GPUs, respectively, demonstrating its efficiency on modern hardware. Finally, to further narrow the gap between compressed and dense models, we use an optional lightweight PEFT method, which provides up to an 1.66%  additional accuracy improvement for LLaMA-2-13B.
@@ 2 Preliminaries
Group quantization (Alistarh et al., 2017; Gunho et al., 2022) tackles AbsMax’s outlier sensitivity by assigning separate scaling factors to subgroups of the weight matrix. This approach captures local variations in weight magnitudes, reducing quantization error for non-uniform distributions. However, storing multiple scaling factors increases memory usage, and subgroup-specific dequantization increase computational complexity, potentially slowing inference. The challenges of using group quantization are discussed in Appendix U.
@@ 3.1 SLiM-Quant Quantization Method
Activation-aware SLiM-Quant. Recent work has shown that the quantization of a subset of the weight channels has a higher impact on output error of the model (Xiao et al., 2023; Lin et al., 2024). We extend SLiM-Quant by incorporating an output error minimization approach inspired by AWQ (Lin et al., 2024). Similar to AWQ, our revised algorithm applies a scaling strategy to activations, reducing the quantization error of salient weight channels. Specifically, we scale up the weights associated with the most significant channels and correspondingly scale down the related input activations. This approach maintains computational equivalence while effectively lowering the quantization-induced output error. In particular, scaling approximately 1% of the channels does not alter the overall quantization parameters but significantly reduces errors in the critical channels.
However, our approach diverges from AWQ by introducing a novel saliency metric that jointly considers both activations and weights. We define the saliency of each channel as the product of the normalized average magnitudes of inputs and weights, expressed as $|diag(\mathbf{x})\times\mathcal{W}|$, where $\mathbf{x}$ and $\mathcal{W}$ denote the average magnitude of activations and weights, respectively, and |.||.|| . | denotes the element-wise absolute value operator. Channels with the highest saliency are scaled by a factor of $s>1$, while their corresponding activations are scaled by $\frac{1}{s}$. Although this method introduces modest computational overhead, that is attributed to on-the-fly adjustments of roughly 1% of activations and resulting irregular memory access patterns, it yields measurable accuracy improvements. These results underscore a clear trade-off between computational complexity and model performance, highlighting the relative strength of SLiM-QuantO (SLiM-Quant with output error minimization) over SLiM-QuantW (SLiM-Quant with weight error minimization). Please note that SLiM-Quant without the superscript W denotes the weight error minimization version of SLiM-Quant .
@@ 3.2 SLiM-LoRA Low-rank Adapters
A straightforward approach minimizes the total error norm between $\mathcal{W}$ and $\mathcal{W}^{C}$, focusing solely on reducing the error magnitude while ignoring the saliency of individual elements in the weight matrix. We call this method Naive-LoRA as it overlooks the importance of individual elements in the weight matrix. However, this method is suboptimal and can be substantially improved.
@@
Assuming that there exists an additive invertible saliency function $F:{\mathbb{R}}^{d_{in}\times d_{out}}\rightarrow{\mathbb{R}}^{d_{in}\times d_{out}}$, we need to solve Equation 8 to find the optimal adapters. Using the additive property of the saliency function $F(.)$, we can simplify Equation 8 to Equation 9.
@@ 4 Experimental Results
Models, Datasets, and Evaluation. [^3] All the experiments in the paper were run at the University of Toronto We evaluate SLiM on the OPT (Zhang et al., 2022) and LLaMA-2 (Touvron et al., 2023) model families, both of which serve as standard baselines in model compression studies (Ma et al., 2024; Frantar & Alistarh, 2023; Sun et al., 2023). Model accuracy is assessed on a range of zero-shot downstream tasks, including MMLU (Hendrycks et al., 2020), Piqa (Bisk et al., 2020), Arc-Easy, Arc-Challenge (Clark et al., 2018), WinoGrande (Sakaguchi et al., 2021), and OpenBookQA (Mihaylov et al., 2018). For zero-shot evaluations, we utilize the Language Model Evaluation Harness (Gao et al., 2024) framework. In line with prior work (Sun et al., 2023; Frantar & Alistarh, 2023; Ma et al., 2024), we also report the perplexity of the models on a language modeling task on the WikiText2 (Merity et al., 2016) dataset, provided in Appendix G.
@@
Baselines. We compare SLiM against state-of-the-art one-shot pruning methods, including Wanda (Sun et al., 2023), SparseGPT (Frantar & Alistarh, 2023), and Magnitude Pruning (Han et al., 2015), as well as one-shot quantization techniques like OPTQ (Frantar et al., 2022), OmniQuant (Shao et al., 2023), AffineQuant (Ma et al., 2024), L2QER (Zhang et al., 2024a), and AbsMax. Additionally, we extend Joint Sparsification and Quantization (JSQ) (Guo et al., 2024) to support 4-bit weight quantization and include it in our experiments. To ensure fairness, we use the optimal hyperparameters reported for each method, or the default hyperparameters if not explicitly reported. For a thorough description of the notations used to show the different variants of SLiM, please see Table 4 in Appendix A. For more details about the hyperparameters used in different experiments, please see Appendix T.
@@
L2QER (Zhang et al., 2024a) and OATS (Zhang & Papyan, 2024) are the two independent and concurrent compression methods utilizing zero-shot low-rank adapters to enhance model accuracy. Our approach, SLiM, significantly diverges from them in several key aspects. First, we employ saliency-based low-rank adapters to mitigate compression loss in quantized and sparse models, whereas L2QER is tailored exclusively for quantization, resulting in reduced accuracy when combined with sparsity, as demonstrated in the subsequent subsections, and OATS is designed for unstructured sparsity only without quantization, which does not have acceleration support on NVIDIA GPUs. Second, we introduce SLiM-Quant , which lowers the overhead and complexity of group quantization compared to methods like L2QER. Finally, SLiM compresses and fine-tunes low-rank adapters efficiently to minimize overhead. In contrast, L2QER and OATS rely on full-precision low-rank adapters, which incur additional overhead and do not benefit from the parameter-efficient fine-tuning proposed in our work.
@@
Naive-LoRA SLiM-QuantW 34.28 33.38 38.36 41.21 44.91 45.25 48.45 51.94
SLiM-LoRA SLiM-QuantW 34.62 34.36 40.61 42.73 45.99 46.09 51.15 54.94
SLiM-LoRAQ SLiM-QuantW 34.43 34.30 40.11 42.37 46.33 46.24 51.02 53.55
@@
Naive-LoRA SLiM-QuantW 34.77 34.23 40.40 43.37 46.64 47.30 51.52 55.33
SLiM-LoRA SLiM-QuantW 35.20 35.32 41.85 43.48 47.08 47.96 54.26 57.85
SLiM-LoRAQ SLiM-QuantW 35.35 35.13 41.74 43.63 47.16 47.86 54.18 57.33
@@
Naive-LoRA + FT SLiM-QuantW 50.89 55.70
SLiM-LoRA + FT SLiM-QuantW 52.12 56.60
SLiM-LoRAQ + FT SLiM-QuantW 48.31 56.50
@@
Naive-LoRA + FT SLiM-QuantW 52.90 57.08
SLiM-LoRA + FT SLiM-QuantW 54.69 57.96
SLiM-LoRAQ + FT SLiM-QuantW 53.57 57.78
Accuracy Results. We evaluate the accuracy of SLiM and other state-of-the-art pruning and quantization methods across 2:4 and unstructured sparsity benchmarks, highlighting SLiM’s superiority in Table 1. SparseGPT and Group OPTQ, designed to work together, achieve competitive performance. For other advanced quantization methods, we pruned models using Wanda and quantized the sparse checkpoints with Group AbsMax, AWQ, OmniQuant, and AffineQuant, reporting the best results (detailed in Appendix H). In particular, methods such as OmniQuant and AffineQuant struggle to quantize OPT-350M, often resulting in NaN values. Moreover, AWQ, OmniQuant, AffineQuant, and L2QER encounter out-of-memory (OOM) errors when compressing models on a single A100-40GB GPU. While JSQ performs well for the LLaMA-2 family, its difficulty compressing the OPT family limits its broader applicability.
Table 2 demonstrates how lightweight fine-tuning (FT) improves the accuracy of both SLiM-LoRA and Naive-LoRA, with SLiM-LoRA exhibiting greater gains due to its saliency-aware design. Further details on the fine-tuning process and its overhead are provided in Appendix K, illustrating its practicality for enhancing compressed model performance.
SLiM with MaskLLM. MaskLLM is the state-of-the-art pruning method designed for 2:4 sparsity. It keeps the original weights in the model intact, while finding the optimal 2:4 masks for the model through a mask training phase. As a result, it can be combined with SLiM to boost the accuracy of the models even further. Table 3 summarizes the average accuracy results of MaskLLM on six zero-shot downstream tasks and its perplexity on WikiText2 dataset.
Table 3: Accuracy (Acc) and perplexity (PPL) of MaskLLM combined with SLiM on LLaMA-2-7B.
Pruning/LoRA Weight LLaMA-2-7B
Method Quantization Acc PPL
Dense - 56.6 5.47
MaskLLM - 49.7 7.3
Naive-LoRA - 52.3 6.6
SLiM-LoRA - 52.2 7.0
Naive-LoRA + FT - 52.6 6.6
SLiM-LoRA + FT - 52.9 6.6
MaskLLM Group AbsMax 49.2 7.6
Naive-LoRA SLiM-Quant 50.5 7.3
SLiM-LoRA SLiM-Quant 51.4 7.5
Naive-LoRA + FT SLiM-Quant 51.2 6.9
SLiM-LoRA + FT SLiM-Quant 52.1 6.8
Large Compressed vs. Small Dense Models. This section compares large compressed models with dense models of equivalent parameter size, offering guidelines for configuration selection under hardware constraints. We focus on 2:4 sparsity due to its hardware acceleration support and evaluate the OPT model family, which spans a wide range of sizes for comprehensive analysis.
We analyze model performance by plotting average accuracy against parameter size, calculated as detailed in Appendix L. This visualization enables a direct performance comparison between models with an equal number of bits.
@@
Speedup. Leveraging sparsity and quantization enhances GPU resource utilization, enabling faster model inference. Following Wanda’s experimental setup, we evaluate the speedup achieved across different model layers and sizes. Similar to Wanda, AWQ, and QuaRot (Ashkboos et al., 2024), we focus on consumer-grade GPUs and conduct our experiments on NVIDIA RTX 3060 GPUs. Speedup results for NVIDIA A100 GPUs are provided in Appendix J.
@@
SLiM achieves notable speedups through optimized sparse and quantized matrix multiplication, utilizing Sparse Marlin (Frantar et al., 2024) integrated with vLLM (Kwon et al., 2023). For inference, we adopt small batch sizes during decoding, as recommended by prior works (Xia et al., 2023; Zheng et al., 2022). Dense Quantized Marlin or PyTorch kernels handle the low-rank adapters based on their quantization status. Figure 3 highlights the speedup achieved across different LLaMA-2 layers compared to dense, unquantized models. The breakdown of the speedup, showing the contribution of the quantization and sparsity, is demonstrated using brighter and darker colors respectively. Larger matrices, such as those in feed-forward modules, consistently yield greater speedups, aligning with trends detailed in Appendix J.
Model Memory Reduction. We evaluate SLiM’s memory reduction on A100 GPUs and our experiments show that SLiMQ achieves 0.23×0.23$\times$0.23 × and 0.23×0.23$\times$0.23 × memory reduction on LLaMA-2-7B and LLaMa-2-13B respectively. The reductions for SLiM are 0.33×0.$3$ and 0.34×0.34$\times$0.34 × respectively.
@@
An evaluation of SLiM with input quantization using FP8 is provided in Input Quantizatoin (Appendix B). The results show that input quantization has a minimal impact on the accuracy of the models using SLiM .
A comparison between weight error minimization and activation error minimization in SLiM-Quant is provided in SLiM-QuantW vs. SLiM-QuantO (Appendix C). The experiments show that the gap between output error minimization and weight error minimization is not significant.
We evaluate SLiM on sparse-only and quantized-only models to isolate their effectiveness. Results in Additional Sparse-only Results (Appendix D) and Additional Quantization-only Results (Appendix E) demonstrate that SLiM and SLiM-Quant consistently outperform state-of-the-art compression methods.
The Language Modeling Experiments (Appendix G) evaluates SLiM across sparse and quantized, sparse-only, and quantized-only models on WikiText-2. The results align with the accuracy trends reported in the main paper, further validating the effectiveness of SLiM .
The Fine-tuning Costs (Appendix K) shows that SLiM reduces fine-tuning overhead from over 36 days for 13B parameter models to just 14 hours on a single GPU, demonstrating its practicality and efficiency.
Figure 3: LLaMA-2 family of models speedup ($\times$) using SLiM compared to original dense unquantized model on NVIDIA RTX-3060. $\uparrow$ shows higher speedup. The brighter color shows the contribution of quantization to the total speedup.
We provide a comparison between Sparsity vs. Quantization (Appendix I) to show that combining 50% sparsity and 4-bit quantization helps achieve better compression results in comparison to solely using 2-bit quantization, while maintaining a similar compression ratio ($\sim$8×8$\times$8 ×).
Additional speedup results for SLiM on NVIDIA A100-40GB GPUs are provided in the Additional Speedup Results (Appendix J). A theoretical analysis of computation and memory reductions can be found in the Computation Reduction Analysis (Appendix M) and Memory Reduction Analysis (Appendix L), highlighting the efficiency of SLiM.
Compression Costs (Appendix N)
details the time required to compress models of various sizes across different methods. Rank Analysis (Appendix O) explores how rank choices in low-rank adapters impact computational and memory costs, as well as model accuracy. Sparsity Analysis (Appendix D) analyzes the effects of different sparsity ratios on model compression. Lastly, Effects of Calibration Sample Count (Appendix P) evaluates the influence of calibration sample counts on the accuracy of calibration-based methods.